OpenStudy (anonymous):

name the axis of symmetry and and the coordinates of vertex y=x^2-4x-3

OpenStudy (anonymous):

Can you complete the square of x^2-4x-3?

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

To complete the square, first divide the coefficient in front of x by 2 and then square it, then add and subtract it in to the expression so (-4/2)^2=4 x^2-4x+4-4-3 the first 3 terms factor as (x-2)^2, combine the remaining constants y=(x-2)^2-7 now its in vertex form y=a(x-h)^2+k where (h,k) is the vertex and x=h is the axis of symmetry

OpenStudy (anonymous):

So vertex is (2,-7) axis of symmetry is x=2

OpenStudy (anonymous):

thanks man i think im getting it now